5kg of potassium is reacted with excess sulfur to make potassium sulfide. Work out the mass of potassium sulfide produced.
1
Expert's answer
2020-10-14T14:03:25-0400

128.2 mole x mole

2K + S -> K2S

2 mole 1 mole


n(K) = m(K)/M(K) = 5 000 g/39 g/mole = 128.2 mole

128.2 mole/2 mole = x mole/1 mole

x = (128.2 * 1)/2 = 64.1 mole = n(K2S)

m(K2S) = n(K2S) * M(K2S) = 64.1 * (39*2 + 32) = 64.1 * 110 = 7051 g = 7.051 g